Operations Report – March 05th

Crew 223 Operations Report 05/03/2020

SOL: 4

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Wednesday generator:
Off at 6:00am
On at 9:15pm SOC was 84%

Thursday generator:
Off at 4:30am
On at 9pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149709.09 units

Water (static tank) – 319 gallons

Static to Loft Pump used – yes

Water in GreenHab – 166.4 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Perseverance (former Demos) rover used: no
Hours: 149.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 133.7
Beginning charge: 100%
Ending charge: 86%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: 1 & 2 used during the EVA. We had trouble starting the
Honda, and finally thought the battery was flat. But during the engineer
EVA, Honda started without any problem (using the starter). So maybe it
was due to the cold weather… Crew members going in EVA tomorrow
morning will try to start it. If non working, we will probably take the
battery in the RAM to charge it. All ATVs were run during the engineer
EVA for battery check.
Oil Added? no
ATV Fuel Used: ~5%
Hours the ATVs were Used today: 20 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? Nothing to report

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 8, 9, 10 were used during the EVA.
Suit 7 is now fixed and perflectly working !

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: The GreenHab officier switched the
working and non working neons of the halogen lamp in order to have more
light on tomatoes than on aromatics.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We successfully fixed suit 7 ! What we found
is a problem on the suit female plug : in a certain position, one of the
cables was in contact with one of the 2 free pins. And thus connecting
the electrical circuit to the ground. Our solution was to isolate those
2 pins with an heat-shrinkable sleeve. Cf the 2 pictures attached. We
did the same thing on suits 8, 9, 10 to prevent them from this.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report